Question 17 Your answer is partially correct. Try again. Culver Company purchased a heavy-duty truck on July 1, 2014, for $31Your answer is partially correct. Try again. Culver Company purchased a heavy-duty truck on July 1, 2014, for $31,320. It was estimated that it would have a useful life of 10 years and then would have a trade-in value of $5,640. The company uses the straight-line method. It was traded on August 1, 2018, for a similar truck costing $46,964; $17,124 was allowed as trade-in value (also fair value) on the old truck and $29,840 was paid in cash. A comparison of expected cash flows for the trucks indicates the exchange lacks commercial substance. What is the entry to record the trade-in? (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)
Answer #1
Please give positive ratings so I can keep answering. If you have any queries please comment. Thanks!
Culver Company
If the transaction has a commercial substance than the transaction are to be recorded as their Fair Value.
Amount $
Truck (cost)     31,320.00 A
Salvage Value       5,640.00 B
Depreciable value     25,680.00 C=A-B
Life (Years)             10.00 D
Annual depreciation       2,568.00 E=C/D
Depreciation for 2014 (6 months)       1,284.00 F=E/12*6 July to Dec 2014 is 6 months.
Depreciation for 2015 (12 months)       2,568.00 G=E
Depreciation for 2016 (12 months)       2,568.00 H=E
Depreciation for 2017 (12 months)       2,568.00 I=E
Depreciation for 2018 (7 months)       1,498.00 J=E/12*7 Jan to July 2018 is 7 months.
Accumulated depreciation     10,486.00 K=F+G+H+I+J
Book value of Truck     20,834.00 L=A-K
Fair value of Truck     17,124.00 M
Cash given up     29,840.00 N
Fair value of Truck exchanged     46,964.00 O
Profit on Truck exchanged      (3,710.00) P=L-M
Journal entry Debit $ Credit $
New Truck     46,964.00
Accumulated depreciation     10,486.00
Old Truck      31,320.00
Cash      29,840.00
Loss on exchange of Truck       3,710.00
    61,160.00      61,160.00